OpenStudy (anonymous):

write in slope intercept form the equation of the line that passes through (1,4) and has the slope 2/1

OpenStudy (yanasidlinskiy):

y=mx+b. This is the slope formula.

OpenStudy (anonymous):

so this gives you y = 2 x + b then 4 = 2(1) + b solve for b